28)One card is drawn from a well shuffled deck of 52 cards.

is equally likely, calculate the probability that the card will be


not a diamond
Answer #1

We assumed in i) above that A is the event card drawn is a diamond .so the event card drawn is not a diamond may be denoted as A or not A 

Now p(not A)=1-P(A)=1-1/4=3/4

answered by: Vijayan